### v1

# Model-independent unbinned analysis of $B\to K^{*}(\to K^{+}\pi^{-})\mu^{+}\mu^{-}$: zeroes, bounds, Wilson coefficients and symmetries

Abstract
We present a model-independent method to study the four-body decay $B\to K^{*}(\to K^{+}\pi^{-})\mu^{+}\mu^{-}$, based on extracting continuous observables with a moments approach. The method allows the observables to be determined unbinned in both the dilepton and $K^{+}\pi^{-}$ invariant masses on which the decay dynamics depend. This will allow the method to shed new light on how the observables depend on the P- and S-wave contributions to the $K^{+}\pi^{-}$ system. This approach contrasts with the state-of-the-art analyses, which bin in dilepton and $K^{+}\pi^{-}$ mass, or use a model for the dependence of the underlying decay amplitudes on these masses. The method does not require making a statistical fit, and so avoids problems of biases and poor uncertainty estimation when dealing with small samples or a large number of fit parameters. We provide the Standard Model predictions for the unbinned optimised observables, derive new geometrical bounds on their values and study the robustness of these bounds in the presence of a scalar new physics contribution. We explore the zero-crossing points of $P_{2}$ and $P_{4,5}^{\prime}$ observables as a function of a new physics contribution to the dominant vector Wilson coefficient, $C_{9}^{\rm NP}$. We also discuss the conditions that can be used to test the theoretical model of the amplitudes needed for an experimental amplitude analysis. Finally, as an illustration, we show how the proposed method might be used to extract the zero-crossing points, make a comparison with the bounds and test a non-trivial relation between the observable values.
I Introduction
Anomalies have been observed in the measurements of a range of $B$-meson decay modes. The most persistent effects are the deviations in the $P_{5}^{\prime}$ Descotes-Genon et al. (2013a) angular observable belonging to the $b\to s\mu^{+}\mu^{-}$ decay $B\to K^{*}(\to K^{+}\pi^{-})\mu^{+}\mu^{-}$, the deficit in the branching fractions of several other $b\to s\ell^{+}\ell^{-}$ processes (with $\ell=e,\mu$Aaij et al. (2014, 2021) and the Lepton Flavour Universality (LFU) ratios $R_{D,D^{*}}$ with charged current $b\to c\tau\nu$ decays Aaij et al. (2023a, b).
The data of the four-body angular distribution $B\to K^{*}(\to K^{+}\pi^{-})\mu^{+}\mu^{-}$ have been commonly presented in bins of the dilepton invariant mass squared (denoted $q^{2}$) by the LHCb collaboration Aaij et al. (2020), but also by the Belle Wehle et al. (2017), CMS Hayrapetyan et al. (2024) and ATLAS Aaboud et al. (2018) collaborations. Analyses are typically made within a window of $K^{+}\pi^{-}$ invariant mass, $m_{K\pi}$. It is possible to construct a full set of angular observables, optimised to have reduced sensitivity to hadronic effects Descotes-Genon et al. (2013b); Capdevila et al. (2018). The LHCb collaboration have recently presented two analyses that determine the complete set of such observables unbinned in $q^{2}$ Aaij et al. (2024a, b).
The purpose of these analyses was to extract Wilson coefficients together with hadronic information directly from the data using an Effective Field Theory (EFT)-based functional form for the transversity amplitudes. These analyses differ in the way that the hadronic contributions are parameterised. On the one hand, Ref. Aaij et al. (2024a) relies on the use of a $z$-expansion to parameterise the non-local charm-loop contributions and includes constraints from $B(B\to K^{*}J/\psi)$ and $B(B\to K^{*}\psi(2s))$ decays. In this analysis, two hypotheses are considered for the data included in the fit: (i) $q^{2}<0$, where constraints coming from the light-cone Operator Product Expansion (OPE) calculation in Ref. Gubernari et al. (2021) are included; and (ii) $q^{2}>0$, where these constraints are excluded. By contrast, the alternative analysis of Ref. Aaij et al. (2024b) uses a dispersive approach based on the resonance model of Ref. Cornella et al. (2020) and data from the decay $B\to K^{*}\mu^{+}\mu^{-}$ alone. An interesting outcome of both of these analyses is that the data indicate that the charm-loop is insufficient to explain the observed anomalies, and a new physics contribution is required.
Our goal in the present paper is to report a new model-independent method to parameterise directly the unbinned observables, as opposed to the amplitudes, without any theoretical input. The results of this method will thus be a facsimile of the experimental data, which can then be used for further analysis of the underlying physics. We also show that this method can be used to test bounds on the observables and to extract zero-crossing points that give additional information on the Wilson coefficients. We use simplified pseudo-experiments to illustrate how the method works. This new method is in contrast with the above analyses, which attempt to parameterise the various contributions to the transversity amplitudes, i.e. local/non-local form factors, resonance lineshapes etc. It also contrasts with another proposed method to directly parameterise the amplitudes with $q^{2}$-dependent polynomials Egede et al. (2015) and the alternative method of Ref. Beck et al. (2025). In all of the above cases, the theoretical models and hypotheses used to construct the amplitudes are then inherited by any subsequent calculation of the observables.
The structure of the paper is as follows. In section II, we present our Standard Model (SM) predictions for the optimised unbinned observables, together with a method to test the validity of any theoretical model of the amplitudes. In section III, we discuss a new method to extract unbinned observables from data, which is underpinned by a moments approach. In section IV, we present some examples of the information that can be extracted from the knowledge of the unbinned shape of the observables. In particular, we present new geometrical bounds on the optimised observables that can be used to test the robustness of an experimental analysis. We explore the sensitivity of the bounds, as well as a relation between the observables in the presence of a scalar. We then discuss the information that can be extracted directly from the zeroes of $P_{2}$ and $P_{4,5}^{\prime}$ and, finally, conclude in section V.
II SM predictions for the unbinned observables
Electroweak transitions of $B$-mesons occur at characteristic energies far below the electroweak scale. This separation of energy scales implies that electroweak degrees of freedom are never resolved in such processes. Consequently, the physics at high energies can be effectively described at low energies within an EFT framework. This effective theory is commonly referred to in the literature as the Weak Effective Theory (WET) Grinstein et al. (1988); Buchalla et al. (1996) or the Low Energy Effective Theory (LEFT) Jenkins et al. (2018). The structure of the corresponding effective Hamiltonian follows an OPE
$\mathcal{H}_{\text{eff}}=-\frac{4G_{F}}{\sqrt{2}}V_{tb}V_{ts}^{*}\sum_{i}\mathcal{C}_{i}(\mu)\mathcal{O}_{i}(\mu),$ (1)
where $G_{F}$ is the Fermi constant, $V_{tb}$ and $V_{ts}$ are CKM matrix elements, $\mathcal{C}_{i}(\mu)$ are the Wilson coefficients encoding the short-distance dynamics, and $\mathcal{O}_{i}(\mu)$ are the corresponding effective operators. The scale $\mu$ represents the renormalisation scale at which the Hamiltonian is evaluated. In addition to the SM operators, the effective Hamiltonian may also include operators that encapsulate structures not generated in the SM, such as right-handed currents or scalar interactions, which arise in various New Physics (NP) scenarios. The most relevant operators for the following discussion are
[…]
where colour indices have been omitted, $P_{L,R}=(1\mp\gamma_{5})/2$ are chirality projection operators, $F^{\mu\nu}\equiv\partial^{\mu}A^{\nu}-\partial^{\nu}A^{\mu}$ is the electromagnetic field strength tensor ($A^{\mu}(x)$ being the photon field), and $\sigma^{\mu\nu}=\frac{i}{2}[\gamma^{\mu},\gamma^{\nu}]$, with $\gamma^{\mu}$ denoting the gamma matrices in four dimensions.
The Wilson coefficients for the most relevant operators of the effective Hamiltonian in Eq. (1) within the SM, evaluated at $\mu_{b}=4.8$GeV, are given in Table 1 Gambino et al. (2003); Bobeth et al. (2004); Misiak and Steinhauser (2007); Huber et al. (2006, 2008).
This paper will be focused on $B$-meson decays into a $K^{*}$ meson and a lepton pair $\ell^{+}\ell^{-}$. The relevant amplitude for this process has the structure
$\mathcal{M}(B\to K^{*}\ell^{+}\ell^{-})=$ $\,\frac{G_{F}\alpha}{\sqrt{2}\pi}V_{tb}V_{ts}^{*}\Big{[}(\mathcal{A}_{V}^{\mu}+\mathcal{H}^{\mu})\bar{u}_{\ell}\gamma_{\mu}v_{\ell}$
$+\mathcal{A}_{A}^{\mu}\bar{u}_{\ell}\gamma_{\mu}\gamma_{5}v_{\ell}\Big{]},$ (3)
with
$\mathcal{A}_{V}^{\mu}=$ $-2m_{b}\dfrac{q_{\nu}}{q^{2}}\,{\cal C}_{7}\langle K^{*}|\bar{s}\sigma^{\mu\nu}P_{R}b|B\rangle+{\cal C}_{9}\langle K^{*}|\bar{s}\gamma^{\mu}P_{L}b|B\rangle,~{}$
$\mathcal{A}_{A}^{\mu}=$ $\,\,{\cal C}_{10}\langle K^{*}|\bar{s}\gamma^{\mu}P_{L}b|B\rangle,~{}$ (4)
and
$~{}\mathcal{H}^{\mu}$ $=\dfrac{-16i\pi^{2}}{q^{2}}\times$ (5)
[…]
where we assume no right-handed currents. The electromagnetic quark current is denoted as $j^{\mu}_{\rm em}$, $q^{\mu}$ represents the dilepton momentum, and $\mathcal{O}_{i=1,\ldots,6,8}$ ($\mathcal{C}_{i=1,\ldots,6,8}$) correspond to four-quark effective operators (Wilson coefficients) in the same effective Hamiltonian.
The matrix elements in Eq. (4) can be expressed in terms of scalar structure functions known as local form factors. Our framework treats the local form factors in the low- and high-$q^{2}$ regions separately. Consequently, it does not attempt or rely on combined fits of theoretical calculations at low (or negative) $q^{2}$ with lattice data at large $q^{2}$.
In the low-$q^{2}$ region, we employ the most recent calculations of the $B\to K^{*}$ form factors within the light-cone OPE. The form factors are computed using light-cone sum rules (LCSRs) with $B$-meson light-cone distribution amplitudes (LCDAs), incorporating terms up to twist four for both two- and three-particle distributions. As described in Ref. Gubernari et al. (2019), these results are then fitted to a $z$-expansion within the so-called BSZ parameterisation Bharucha et al. (2016). We use the resulting $z$-expansion parameterisation of the form factors and decompose them into three pieces: i) soft form factors $\xi_{\perp,\parallel}$ Charles et al. (1999); Beneke and Feldmann (2001), which are the relevant structure functions emerging in the formal Heavy Quark Effective Theory (HQET) and Soft Collinear Effective Theory (SCET) limits ($m_{b}\to\infty$ and $q^{2}\to 0$); ii) $O(\alpha_{s})$ corrections $\Delta F^{\alpha_{s}}$, which are calculable within QCD factorization (QCDf) Beneke and Feldmann (2001); and iii) factorisable power corrections $\Delta F^{\Lambda}$ Descotes-Genon et al. (2014); Capdevila et al. (2017),
$F(q^{2})=F^{\infty}(\xi_{\perp}(q^{2}),\xi_{\|}(q^{2}))+\Delta F^{\alpha_{s}}(q^{2})+\Delta F^{\Lambda}(q^{2}).$ (6)
The latter, representing part of the $O(\Lambda_{\rm QCD}/m_{b})$ power corrections to QCDf, is parameterised as an expansion in $q^{2}/m_{B}^{2}$ Jäger and Martin Camalich (2013),
$\Delta F^{\Lambda}(q^{2})=a_{F}+b_{F}\,\frac{q^{2}}{m_{B}^{2}}+c_{F}\,\frac{q^{4}}{m_{B}^{4}}+\ldots\;,$ (7)
where $F=V,A_{0,1,2},T_{1,2,3}$. For each form factor, the parameters $a_{F}$, $b_{F}$, and $c_{F}$ are determined from a fit to the form factor values, taking into account the corresponding structure in the HQET/SCET limit, as well as the associated $O(\alpha_{s})$ corrections, following the schematic structure implied in Eq. (6). Further details on the fit strategy and results for the factorisable power corrections, as extracted from KMPW Khodjamirian et al. (2010) and BSZ Bharucha et al. (2016) form factors, can be found in Refs. Descotes-Genon et al. (2014); Capdevila et al. (2017).
In contrast, in the high-$q^{2}$ region we use lattice QCD calculations of $B\to K^{*}$ form factors Horgan et al. (2014), which are subsequently fitted using a $z$-expansion within the BCL parameterisation Bourrely et al. (2009).
Model ${\cal C}_{7}^{\rm eff}(\mu_{b})$ ${\cal C}_{9}^{\rm eff}(\mu_{b})-Y(q^{2})$ ${\cal C}_{10}(\mu_{b})$
SM -0.29 +4.08 -4.31
Model ${\cal C}_{7}^{NP}$ ${\cal C}_{9}^{U}$ ${\cal C}_{10}^{U}$ ${\cal C}_{7^{\prime}}$ ${\cal C}_{9^{\prime}}^{U}$ ${\cal C}_{10^{\prime}}^{U}$
NP1 -1.17
NP2 +0.00 -1.21 +0.07 +0.01 -0.04 -0.06
Table 1: Main SM and NP Wilson coefficients at a scale $\mu_{b}=4.8$ GeV. The remaining SM Wilson coefficients can be found in Ref. Descotes-Genon et al. (2013b). NP1 hypothesis corresponds to the preferred LFU scenario discussed in Ref. Algueró et al. (2023), while NP2 corresponds to the 6D LFU hypothesis.
While the contributions from local form factors are included in $\mathcal{A}_{V,A}^{\mu}$, non-local charm-quark loop contributions appear in the term $\mathcal{H}^{\mu}$. These effects are accounted for by recasting the phenomenological parameterisation provided in Eq. (7.14) of Ref. Khodjamirian et al. (2010). This parameterisation is itself based on the light-cone OPE and LCSR calculations of the non-local form factor in the region $q^{2}\ll 4m_{c}^{2}$, and its corresponding matching to the dispersion relation proposed in the same paper. In practice, this is implemented as a shift in the factorisable charm-quark loop function $Y(q^{2})$ Buras et al. (1994)
$Y_{i}(q^{2})=Y(q^{2})+s_{i}Y_{i}^{LD}(q^{2}),\quad i=\perp,\parallel,0,$ (8)
where the long-distance functions $Y_{i}^{LD}$, which depend on the $K^{*}$ polarization states $\perp$, $\parallel$, and 00, are given by Descotes-Genon et al. (2016)
$Y_{LD}^{\perp}(q^{2})$ $=\dfrac{a^{\perp}+b^{\perp}q^{2}(c^{\perp}-q^{2})}{q^{2}(c^{\perp}-q^{2})},$ (9)
$Y_{LD}^{\parallel}(q^{2})$ $=\dfrac{a^{||}+b^{||}q^{2}(c^{||}-q^{2})}{q^{2}(c^{||}-q^{2})},$ (10)
and similarly for the longitudinal amplitude, which contains no pole at $q^{2}=0$
$Y_{LD}^{0}(q^{2})=\dfrac{a^{0}+b^{0}(q^{2}+s_{0})(c^{0}-q^{2})}{(q^{2}+s_{0})(c^{0}-q^{2})},$ (11)
where $s_{0}=1$ GeV2. The parameters $a^{\perp,\parallel,0}$, $b^{\perp,\parallel,0}$, and $c^{\perp,\parallel,0}$ are determined by ensuring that they reproduce the phenomenological model of Ref. Khodjamirian et al. (2010) in the $q^{2}$ range from 1 $1\text{ GeV}^{2}$ to 9 $9\text{ GeV}^{2}$. Specific values for these coefficients can be found in Ref. Descotes-Genon et al. (2016).
In order to be conservative, the $s_{i}$ parameters, which describe the interference between long and short distance in Eq. (8), are varied independently for each amplitude within the range $[-1,1]$.
II.1 Definition of unbinned observables
Using the previous framework we compute in this subsection the SM predictions for the unbinned optimised observables.
The definitions of the unbinned observables (counterparts of the binned ones defined in Ref. Descotes-Genon et al. (2013b)) are given by
$P_{1}(q^{2})$ $=\frac{1}{2}\frac{{}[J_{3}+\bar{J}_{3}]}{{}[J_{2s}+\bar{J}_{2s}]}\ ,$ $P_{1}^{\rm CP}(q^{2})$ $=\frac{1}{2}\frac{{}[J_{3}-\bar{J}_{3}]}{{}[J_{2s}+\bar{J}_{2s}]}\ ,$
$P_{2}(q^{2})$ $=\frac{1}{8}\frac{[J_{6s}+\bar{J}_{6s}]}{[J_{2s}+\bar{J}_{2s}]}\ ,$ $P_{2}^{\rm CP}(q^{2})$ $=\frac{1}{8}\frac{[J_{6s}-\bar{J}_{6s}]}{[J_{2s}+\bar{J}_{2s}]}\ ,$
${P_{3}}(q^{2})$ $=-\frac{1}{4}\frac{[J_{9}+\bar{J}_{9}]}{[J_{2s}+\bar{J}_{2s}]}\ ,$ $P_{3}^{\rm CP}(q^{2})$ $=-\frac{1}{4}\frac{[J_{9}-\bar{J}_{9}]}{[J_{2s}+\bar{J}_{2s}]}\ ,$
$P^{\prime}_{4}(q^{2})$ $=\frac{1}{{\cal N}^{\prime}}[J_{4}+\bar{J}_{4}]\ ,$ ${P^{\prime}_{4}}^{\rm CP}(q^{2})$ $=\frac{1}{{\cal N}^{\prime}}[J_{4}-\bar{J}_{4}]\ ,$
$P^{\prime}_{5}(q^{2})$ $=\frac{1}{2{\cal N}^{\prime}}[J_{5}+\bar{J}_{5}]\ ,$ ${P^{\prime}_{5}}^{\rm CP}(q^{2})$ $=\frac{1}{2{\cal N}^{\prime}}[J_{5}-\bar{J}_{5}]\ ,$
${P^{\prime}_{6}}(q^{2})$ $=\frac{-1}{2{\cal N}^{\prime}}[J_{7}+\bar{J}_{7}]\ ,$ ${P^{\prime}_{6}}^{\rm CP}(q^{2})$ $=\frac{-1}{2{\cal N}^{\prime}}[J_{7}-\bar{J}_{7}]\ ,$
${P^{\prime}_{8}}(q^{2})$ $=\frac{-1}{{\cal N}^{\prime}}[J_{8}+\bar{J}_{8}]\ ,$ ${P^{\prime}_{8}}^{\rm CP}(q^{2})$ $=\frac{-1}{{\cal N}^{\prime}}[J_{8}-\bar{J}_{8}],$ (12)
together with
$F_{L}=-\frac{[J_{2c}+\bar{J}_{2c}]}{[d\Gamma/dq^{2}+d\bar{\Gamma}/dq^{2}]}$ (13)
and where the normalization ${\cal N}^{\prime}$ is defined as
${\cal N}^{\prime}=\sqrt{-[J_{2s}+\bar{J}_{2s}][J_{2c}+\bar{J}_{2c}]}\ .$ (14)
The uncertainties for the SM predictions are computed using a Monte Carlo sampling approach, which assumes a multivariate Gaussian distribution for the nuisance parameters described earlier, along with variations in the renormalisation scale, masses, and other relevant factors. The angular observables gives rise to non-linear behaviour, particularly stemming from the normalisation terms. A Gaussian scan of the parameters can therefore lead to non-Gaussian distributions of the observables. This is particularly the case when the normalisation is close to zero. To extract meaningful uncertainties we apply a rescaling factor of $1/3$ to linearise the uncertainty propagation.
Figure 1: SM prediction for the unbinned $F_{L}$ and $P_{1,2,3}$ optimised observables.
Figure 2: SM prediction for the unbinned $P_{4,5,6,8}^{\prime}$ optimised observables.
The SM prediction for the optimised observables $P_{i}$ and $F_{L}$ are shown graphically in Figs. 1 and 2. The $P_{i}^{\rm CP}$ are, to a high accuracy, very close to zero and we do not include them here. For completeness we also provide the non-optimised observables $S_{i}$ and the branching fraction in Appendix A. The unbinned definitions of the $S_{i}$ observables can be easily derived following the same procedure as in Eq. 12. Some remarks are in order here:
  1. i)
    We have explicitly checked that binning the Monte Carlo sampling recovers the binned observable predictions presented in Ref. Algueró et al. (2023);
  2. ii)
    Predictions are shown in the $[0.1,8.0]\,\,\text{GeV}^{2}$ (light quark resonances are neglected) and $[15.0,q^{2}_{\rm max}]\,\,\text{GeV}^{2}$ regions in order to avoid the $c\bar{c}$ sharp resonance regions, which cannot be described by short distance physics. In the high-$q^{2}$ region, we assume quark hadron duality Lyon and Zwicky (2014), due to the large number of broad resonances. However, particular caution needs to be taken in this region, as an unbinned analysis would capture the details of the resonances and thus quark hadron duality violations;
  3. iii)
    The particular behaviour of the $P_{2}$ observable at low-$q^{2}$ is due to a convolution of the geometrical bound ($P_{2}\lesssim 1/2$) described in Eq. (44) below and the asymmetric uncertainty;
  4. iv)
    The observables $P_{3}$ and $P_{6,8}^{\prime}$ are determined by the imaginary components of specific combinations of the helicity amplitudes, making them particularly sensitive to strong phases. The sudden increase in uncertainty seen in Figs. 1 and 2 at high $q^{2}$ comes from the opening of the $c\bar{c}$ threshold, which introduces significant strong phases into $\mathcal{C}_{9}^{\text{eff}}$.
The next issue to address is how to approach unbinned data. Here there are two main options: i) either one parameterises the amplitudes and computes the observables from these amplitudes or ii) one determines the observables directly. In the former case, the observables depend on the model used to parameterise the amplitudes. For example, whether the contribution from NP scalars, tensors or additional operators beyond the SM are included, the modelling of hadronic contributions and the treatment of $m_{K\pi}$ and $q^{2}$. There are different analyses in the literature parameterising the amplitudes Aaij et al. (2024a, b). Instead, the latter case, determining observables, is model-independent. In this way we extract the information without hypothesis a priori and we can perform several tests on the self-consistency of the analysis. In the following subsections II.2 and II.3 we define the tests that a consistent theoretical model of the amplitudes should fulfill. In section III, we describe a model-independent method to determine unbinned observables using a moments approach.
II.2 Testing a theoretical model of the amplitudes: set of conditions
A parameterisation of the amplitudes using an insufficient number of degrees of freedom or with some of these degrees of freedom fixed to values inconsistent with the data will naturally lead to tension with the data.
One possible method to decide, in a systematic way, if a theoretical model for the amplitudes gives a good description of data is if it fulfills a set of conditions. To define these conditions, assuming a $K\pi$ system in a P-wave configuration, it is instructive to introduce a set of three complex 2-component vectors
$n_{\|}=\binom{A_{\|}^{L}}{A_{\|}^{R*}}\ ,\quad\!n_{\bot}=\binom{A_{\bot}^{L}}{-A_{\bot}^{R*}}\ ,\quad\!n_{0}=\binom{A_{0}^{L}}{A_{0}^{R*}}\ ,\quad\!$ (15)
that collect the L and R components of the amplitudes. If one includes the S-wave contribution, two more vectors are needed (see Ref. Algueró et al. (2021)) with corresponding additional conditions.
If a parameterisation of the amplitudes does not respect the conditions, for instance, by fixing some degrees of freedom with theory in a way that is inconsistent with the data, this can manifest itself as the inability to reproduce one or more observables. In order to remedy this, more degrees of freedom must be included or some of the fixed degrees of freedom must be allowed to vary.
In the case where the mass of the leptons is neglected (referred to as the “massless case”), with no scalar NP contribution, the number of degrees of freedom in a P-wave $B\to K^{+}\pi^{-}\mu^{+}\mu^{-}$ decay is eight (see Ref. Egede et al. (2010) and more detailed discussion in subsection IV.1). In the case where the lepton mass is not neglected (referred to as the “massive case"), again with no scalar contribution, there are ten degrees of freedom. Consequently, the number of conditions to test a theoretical model for the amplitudes in the massive case is ten. In the following, we provide the five conditions (three of them complex) that should be fulfilled
${\mathrm{Cond.}}_{1}:\quad|n_{\perp}|^{2}$ $=$ $\frac{2}{\beta^{2}}J_{2s}+\frac{1}{\beta^{2}}J_{3},$
${\mathrm{Cond.}}_{2}:\quad|n_{\|}|^{2}$ $=$ $\frac{2}{\beta^{2}}J_{2s}-\frac{1}{\beta^{2}}J_{3},$
${\mathrm{Cond.}}_{3}{\rm(Real\,and\,Imag)}:\quad n_{\|}^{\dagger}n_{\perp}$ $=$ $\frac{1}{2\beta}J_{6s}{+}i\frac{1}{\beta^{2}}J_{9},$
${\mathrm{Cond.}}_{4}{\rm(Real\,and\,Imag)}:\quad n_{\|}^{\dagger}n_{0}$ $=$ $\frac{\sqrt{2}}{\beta^{2}}J_{4}{+}i\frac{1}{\beta\sqrt{2}}J_{7},$
${\mathrm{Cond.}}_{5}{\rm(Real\,and\,Imag)}:\quad n_{\perp}^{\dagger}n_{0}$ $=$ $\frac{1}{\beta\sqrt{2}}J_{5}{+}i\frac{\sqrt{2}}{\beta^{2}}J_{8},$
where $\beta=\sqrt{1-4m_{\ell}^{2}/q^{2}}$ and $m_{\ell}$ is the lepton mass. These conditions are common to the massless case, where $\beta=1$.
The two extra conditions in the massive case (conditions 6 and 7) that complete the system up to ten equations cannot be written in terms of $n_{i}$ and are given by
$4\frac{m_{\ell}^{2}}{q^{2}}{\rm Re}(A_{\perp}^{L}A_{\perp}^{R*}+A_{\|}^{L}A_{\|}^{R*})=J_{1s}-\left(\frac{2+\beta_{\ell}^{2}}{\beta_{\ell}^{2}}\right)J_{2s},$ (17)
$4\frac{m_{\ell}^{2}}{q^{2}}\left({|A_{t}|^{2}+2\rm Re}(A_{0}^{L}A_{0}^{R*})\right)=J_{1c}+\frac{1}{\beta_{\ell}^{2}}J_{2c}.\quad\quad\quad$ (18)
In the presence of scalar NP, the last two conditions in Eq. LABEL:nJ are affected and also the last condition in Eq. 17. In both cases, the left hand side (LHS) of these conditions has to be modified. In a similar way, one can write a corresponding set of equations for the CP conjugate $J$ and $n_{i}$.
Notice also that the condition
$|n_{0}|^{2}=-\frac{1}{\beta^{2}}J_{2c},$ (19)
is not included, because $n_{0}$ is not independent. Namely, the vector $n_{0}$ can be written in terms of $n_{\|}$ and $n_{\perp}$ Algueró et al. (2021)
$n_{0}=\alpha_{0}n_{\|}+\beta_{0}n_{\perp},$ (20)
where
$\alpha_{0}$ $=$ $\frac{|n_{\perp}|^{2}(n_{\|}^{\dagger}n_{0})-(n_{\|}^{\dagger}n_{\perp})(n_{\perp}^{\dagger}n_{0})}{|n_{\|}|^{2}|n_{\perp}|^{2}-|n_{\perp}^{\dagger}n_{\|}|^{2}},$
$\beta_{0}$ $=$ $\frac{|n_{\|}|^{2}(n_{\perp}^{\dagger}n_{0})-(n_{\perp}^{\dagger}n_{\|})(n_{\|}^{\dagger}n_{0})}{|n_{\|}|^{2}|n_{\perp}|^{2}-|n_{\perp}^{\dagger}n_{\|}|^{2}}.$ (21)
II.3 Interpretation of conditions
In order to test a given amplitude parameterisation using the conditions in Eq. II.2 and Eq. 17, on the LHS of these expressions one should insert the parameterisation of the amplitudes as a function of the relevant free parameters (phases and moduli); and on the RHS, experimental values for the coefficients $J$. A fit for the free parameters can then be used to check if it is possible to correctly reproduce the RHS of each equation.
For illustrative purposes, an approximation to this test  [^1] This is an approximation as lattice information is fitted together with the experimental data. can be implemented by combining the two fits presented in Ref. Aaij et al. (2024a):
  • Fit I:
    A fit to data using only $q^{2}>0$ information but adding lattice calculations. The fit for the observables in Ref. Aaij et al. (2024a) reproduces the binned data very well. We therefore take this fit for the observables as the “data”, i.e., the RHS of the conditions in Eq. II.2 and Eq. 17. The results of this fit are depicted in blue in Fig. 3 and in Fig. 15 in Appendix. B.
  • Fit II:
    A fit to data using, in addition, the $q^{2}<0$ information. This fit is mainly dominated by theory and we reinterpret it as the LHS of the conditions. The results of this fit are depicted in red in Fig. 3 and in Fig. 15.
[…]
This discrepancy indicates a problem in the parameterisation of ${\rm Im}(n_{\|}^{\dagger}n_{0})$. Several reasons can be behind this problem: a) the strong phases fixed using theory at $q^{2}<0$ are in tension with the $B\to K^{*}(\to K^{+}\pi^{-})\mu^{+}\mu^{-}$ data, or higher-order corrections to them are needed; there is a problem in b) the extrapolation to positive $q^{2}$ or c) the lattice calculation. However b) and c) are less plausible, given that the extrapolation and the lattice work well in the case of all the other conditions.
The problem observed in condition 4 is linked to the problem observed in the comparison between ${\rm Im}A_{\|}$ for $q^{2}>0$ and $q^{2}<0$ in Ref. Aaij et al. (2024a). To a much lesser extent there is also a problem in ${\rm Im}A_{\perp}$, as discussed below.
It is illustrative to observe that if one uses the explicit form of the amplitude, as given in Ref. Aaij et al. (2024a):
${\cal A}_{\lambda}^{L,R}$ $=$ ${\cal N}_{\lambda}\left(\left[C_{9}\mp C_{10}\right]{\cal F}_{\lambda}(q^{2})+\frac{2m_{b}M_{B}}{q^{2}}\left[{\cal F}_{\lambda}^{T}(q^{2})\right.\right.$ (22)
$\left.\left.-16\pi^{2}\frac{M_{B}}{m_{b}}{\cal H}_{\lambda}]\right)\right)$
where ${\cal N}_{\lambda}$ stands for the helicity dependent normalisation factor, ${\cal F}_{\lambda}$ and ${\cal F}^{T}_{\lambda}$ are the local form factors and ${\cal H}_{\lambda}$ the non-local correlator. The expression above implies that ${\rm Im}A_{i}^{L}={\rm Im}A_{i}^{R}$ (with $i=\perp,\|,0$). If there is a problem with ${\rm Im}A_{\|}$ it should be more clearly manifest in $P_{6}^{\prime}$. The reason is that this observable is proportional to $J_{7}$, which has the following structure
$J_{7}=\sqrt{2}\beta({\rm Im}A_{\|}^{L}({\rm Re}A_{0}^{R}-{\rm Re}A_{0}^{L})+{\rm Im}A_{0}^{L}({\rm Re}A_{\|}^{L}-{\rm Re}A_{\|}^{R})).$ (23)
Using Eq.22 this implies
$J_{7}=2\sqrt{2}\beta C_{10}\left[({\cal F}_{0}{\cal N}_{0}){\rm Im}A_{\|}^{L}-({\cal F}_{\|}{\cal N}_{\|}){\rm Im}A_{0}^{L}\right].$ (24)
As this is a difference of two small quantities, it is particularly sensitive to their precision. Another coefficient of the angular distribution that is sensitive to ${\rm Im}A_{\|}^{L}$ is $J_{9}$. However in this case,
$J_{9}=\beta^{2}\left[{\rm Im}A_{\perp}^{L}({\rm Re}A_{\|}^{L}+{\rm Re}A_{\|}^{R})-{\rm Im}A_{\|}^{L}({\rm Re}A_{\perp}^{L}+{\rm Re}A_{\perp}^{R})\right].$ (25)
Given that ${\rm Re}A_{\perp}$ and ${\rm Re}A_{\|}$ have opposite signs, the $J_{9}$ coefficient is then not a difference between two small quantities as before, and is consequently less sensitive to the precision in the imaginary parts than $J_{7}$. A marginal tension is still observed in the imaginary component of condition 3 at very low $q^{2}$ and in the imaginary component of condition 5 above 3 GeV2, due to ${\rm Im}A_{\perp}$.
On the contrary we should not expect any sensitivity to ${{\rm Im}A_{\|}}^{L}$ and ${{\rm Im}A_{\|}}^{L}$ in the $P_{2}$ and $P_{5}^{\prime}$ observables, as the dependence on the imaginary parts cancels exactly for the two $J$’s involved in the numerator of these observables
$J_{5}$ $=$ $\sqrt{2}\beta({\rm Re}A_{0}^{L}{\rm Re}A_{\perp}^{L}-{\rm Re}A_{0}^{R}{\rm Re}A_{\perp}^{R}),$
$J_{6s}$ $=$ $2\beta({\rm Re}A_{\|}^{L}{\rm Re}A_{\perp}^{L}-{\rm Re}A_{\|}^{R}{\rm Re}A_{\perp}^{R}).$ (26)
In the case of $P_{4}^{\prime}$, the dependence on ${\rm Im}A_{\|}^{L}$ via $J_{4}$ is hidden by the dominant real parts,
$\!\!J_{4}=\frac{\beta^{2}}{\sqrt{2}}(2{\rm Im}A_{0}^{L}{\rm Im}A_{\|}^{L}+{\rm Re}A_{0}^{L}{\rm Re}A_{\|}^{L}+{\rm Re}A_{0}^{R}{\rm Re}A_{\|}^{R}).$ (27)
In all other $J_{i}$, the real parts dominate over the imaginary ones.
A possible way to fix the problem in condition 4 could be to leave one strong phase free in the amplitude $A_{\|}$ in Fit II and let the observables (in particular $P_{6}^{\prime}$) fix the value of this phase. One should then understand how to obtain this value for the phase from a direct theoretical computation. The same thing could be done for $A_{\perp}$.
Figure 3: The upper panels display the real and imaginary parts of the “condition” $\mathrm{Cond.}_{4}$ in Eq. (II.2), computed from generated samples of the amplitudes following the distribution of fit results from the LHCb analysis of Ref. Aaij et al. (2024a). We present results for the two variants of the analysis described in the text. The lower panel shows the difference between the two curves in the upper panel.
III A new method to determine unbinned observables: A “continuous” moments method
Previous measurements of the angular observables of the $B\to K^{*}\mu^{+}\mu^{-}$ decay have binned the experimental data in $q^{2}$ and integrated the observables over the analysed window of $K^{+}\pi^{-}$ invariant mass, $m_{K\pi}$. These measurements include performing a fit of the three-dimensional angular PDF, with the observables as fit parameters Aaij et al. (2020); Hayrapetyan et al. (2024), or calculations of the corresponding angular moments in each bin Beaujean et al. (2015a); Aaij et al. (2016a, b). By obtaining the complete basis of angular observables, the results are readily interpretable by third parties and may be averaged and re-interpreted in perpetuity. However, by binning the data in $q^{2}$, some information about the underlying physics is invariably lost. Thus far, attempts to exploit the full $q^{2}$ dependence of the decay have involved trying to extract the theory parameters (Wilson coefficients, local and non-local form factor parameters etc.) directly Aaij et al. (2024c, a, b). These measurements are model-dependent, which precludes easy re-interpretation with new theory insights or combination with other experimental data. Furthermore, the $m_{K\pi}$ dependence of the angular observables has not been studied and so experimental results have been interpreted with the approximation that the $K^{\ast{}0}$ is a narrow resonance.
Here we introduce a novel method to extract the full set of model-independent angular observables continuously in $q^{2}$ and $m_{K\pi}$. We do this by extending the previous binned method-of-moments technique into an unbinned one. This is achieved with the use of a $k$-nearest neighbours (knn) algorithm Cover and Hart (1967). The knn with uniform weighting is exactly the definition of the normalised moment (the moment divided by the total weight), $M_{i}$, but instead of being binned, it is really a reflection of the density of the data in the locale of a given phase-space point. Thus one can pick an arbitrary point in $q^{2}$, $m_{K\pi}$, or $(q^{2},m_{K\pi})$, and estimate the value of the angular moment there.
The result of the $k$-nearest neighbours algorithm may be augmented with further processing, for example, Gaussian process regression Pedregosa et al. (2011) or a Savitzky-Golay filter Savitzky and Golay (1964) (other filtering algorithms are available), to produce smoothly varying, continuous, numerical functions. For the studies presented here, a Savitzky-Golay filter is used with polynomial order four. The variances of the observables are calculated with the second-order moments, for which the knn algorithm is again used with subsequent filtering. Covariances (and by implication correlations) may similarly be obtained with the second-order moments. The validity of the second-order moments as an estimate of the statistical uncertainty of the observables is checked with pseudo-experiments by comparing the range of values obtained when sampling with replacement, with the expected range from the calculated variance. The agreement between the two is good for all observables. An illustration of the determination of the observable $S_{5}$ using this method is shown in Fig. 4, resulting from a pseudo-experiment described below.
With this unbinned method, different points in the phase space are statistically correlated with each other. The degree of correlation depends on the number of neighbours in the knn, the smoothing and how far apart the points are. When using the results of this method in further analysis for the underlying physics parameters, knowing these correlations is vital. Consequently, if the correlations are well understood, the precise choice of the number of neighbours or smoothing strength is to some extent a matter of aesthetics. Few neighbours leads to a large apparent uncertainty at a single point but nearby points are relatively independent and so give equal statistical power. Conversely, a large number of neighbours reduces the apparent uncertainty but leads to large statistical correlations with neighbouring regions. What really matters is choosing the analysis parameters that maximise the sensitivity to what is being measured, and in particular allows for the resolution of the phase-space dependence of the observables. Care must be taken, as using too many neighbours can lead to a poor description of the true variables in regions where they vary quickly. However, for some analyses, such as finding the zero-crossing points in $q^{2}$ (discussed in the following sections), one may choose a larger number of neighbours, as the zeros lie in a $q^{2}$ region where the observables of interest are expected to vary smoothly.
A key benefit of an approach using moments is that the observables are obtained with calculations, rather than via a statistical likelihood fit. With a moments calculation one need not be concerned about fits that do not converge, poor statistical behaviour and biased likelihood estimators. Dealing with these has been a major factor in the previous generations of experimental analyses. With this unbinned method one can improve the $q^{2}$ resolution of the angular observables, or investigate the $m_{K\pi}$ dependence of the observables, without having to resort to fits in narrow bins with limited sample sizes, which would involve significant statistical complexity.
Figure 4: Extraction of (top) $S_{5}$ and (bottom) $P_{5}^{\prime}$ as a function of $q^{2}$ from two pseudo-experiments. The bands show the obtained ±$\pm 1\,\sigma$ uncertainty bounds for each pseudo-experiment. The solid lines are the predictions from which the respective data sets were generated: (blue) the SM and (orange) NP1, both from Table 1.
Previous experimental analyses have assumed that the leptons are massless, reducing the number of angular terms to be measured and so simplifying the task. However, it is found that the effects of lepton mass in the muon mode are non-negligible even at relatively large values of $q^{2}$ ($\sim 4\,\,\text{GeV}^{2}$), as shown in Fig. 5. In the absence of S-wave contributions, the full set of angular observables may be extracted with the unbinned moments, fully accounting for the effects of the lepton masses. The observables $S_{1}^{s}$, $S_{1}^{c}$, $S_{2}^{s}$, $S_{2}^{c}$ do not have orthogonal angular functions and must be extracted from a linear combination of the corresponding moments
$S_{1}^{s}$ $=\frac{1}{16}\left(21M_{1}^{s}-14M_{1}^{c}+15M_{2}^{s}-10M_{2}^{c}\right),$
$S_{1}^{c}$ $=\frac{1}{8}\left(-7M_{1}^{s}+28M_{1}^{c}-5M_{2}^{s}+20M_{2}^{c}\right),$
$S_{2}^{s}$ $=\frac{1}{16}\left(15M_{1}^{s}-10M_{1}^{c}+45M_{2}^{s}-30M_{2}^{c}\right),$
$S_{2}^{c}$ $=\frac{1}{8}\left(-5M_{1}^{s}+20M_{1}^{c}-15M_{2}^{s}+60M_{2}^{c}\right).$ (28)
Figure 5: Comparison of $S_{1}^{c}$ and $-S_{2}^{c}$ for the SM prediction from Table 1. The solid lines show the predictions and should align if the effects of the lepton mass are negligible. The band shows the extraction of these observables with the unbinned moments when assuming massless leptons.
As always, contamination from S-wave decays to $K^{+}\pi^{-}\mu^{+}\mu^{-}$ introduces further complications. The first is that the normalised moments provided by the $knn$ algorithm have a factor of $\Gamma_{P}+\Gamma_{S}$ in the denominator, implying that the resulting observables are normalised by $\Gamma_{P}+\Gamma_{S}$, denoted by $\tilde{S}_{i}$ for the P-wave, rather than the usual normalisation by the P-wave only rate, $\Gamma_{P}$, that defines the $S_{i}$ observables. This may be readily rectified by multiplying the P-wave $\tilde{S_{i}}$ observables by a factor of $1/(1-2\tilde{S}_{1a}^{c}+\frac{2}{3}\tilde{S}_{2a}^{c})$ to recover the appropriate normalisation.
The second complication is that there is no single unique linear combination of moments that can be used to extract the complete set of angular observables if lepton masses are considered. The problem arises from terms such as $S_{1}^{s}\sin^{2}\theta_{K}+S_{1}^{c}\cos^{2}\theta_{K}+\tilde{S}_{1a}^{c}$; here one can scale $S_{1}^{s}$ and $S_{1}^{c}$ to accommodate changes in $\tilde{S}_{1a}^{c}$, giving rise to a degeneracy in the determination of these observables. The same problem applies for $S_{2}^{s}$, $S_{2}^{c}$ and $\tilde{S}_{2a}^{c}$.
There are two possible ways to handle this degeneracy. The first is to make approximations and introduce relations between the angular observables, as is done for the current generation of binned analyses. For $q^{2}>4\,\text{GeV}^{2}$ such an approach introduces negligible bias. For lower $q^{2}$ values, one need not utilise all the relations pertaining to negligible lepton mass, instead choosing only those that result in negligible bias in the extracted observables. The second option is to use the information from the variable $m_{K\pi}$ to estimate the relative S- and P-wave contributions. This can be done as part of the moments calculation, assuming some analytic forms for the lineshapes of the P- and S-wave contributions. A full discussion of dealing with S-wave contamination is contained in Appendix D along with a simple study. Unless explicitly stated, the studies presented here only involve the P-wave component of the decay.
Pseudo-data are generated and analysed to prove the principle of the new method. The signal yield from the LHCb experiment with $50\,{\rm fb^{-1}}$ of $pp$ collision data (expected by the end of LHCb upgrade I Aaij et al. (2024d)), across the full $q^{2}$ range, is estimated to be 130,000 candidates and so each pseudo-data set is generated with this quantity of signal. The angular distributions generated follow the SM and NP predictions in Table 1. Here, backgrounds and the warping of the distributions of the kinematic variables due to detector acceptance and selection effects are neglected. In an analysis of collision data the treatment of such effects could follow that of previous moments analyses, such as in Ref. Aaij et al. (2016a). Although, the yields mentioned here pertain to LHCb, the method could readily be used to analyse data from Belle II or the general purpose LHC detectors.
The optimised angular observables, $P_{i}$ and $P_{i}^{\prime}$, are not directly related to the moments but may be calculated from the extracted continuous functions of the non-optimised observables. As the optimised observables are ratios of the $S_{i}$, the normalising rate terms cancel and the $\tilde{S}_{i}$ observables may be used for such a calculation. Propagating the covariance of the moments to the optimised basis of observables is also straightforward. An example of the result of this method is shown in Fig. 4 for $S_{5}$ and the corresponding optimised observable $P_{5}^{\prime}$. The method may be readily extended to assess the angular $CP$-asymmetries by flipping the sign of the moment that contributes to the sum in the $knn$ for $B^{0}$ flavour decays, relative to $\bar{B}^{0}$.
The method has the ability to extract the dependence of the observables on $m_{K\pi}$, or to consider $q^{2}$ and $m_{K\pi}$ together. Such a result is demonstrated in Fig. 6. By considering $m_{K\pi}$, one may readily discern if the finite width of the $K^{\ast{}0}$ leads to a variation of the angular distribution across the width of the $K^{\ast{}0}$. Additionally, by extracting the S-wave moments one is able for the first time to model-independently ascertain the lineshape of the S-wave component of the data. This is particularly important as it is not obvious that the various S-wave resonances (and the non-resonant S-wave contribution) should all have the same angular distributions.
Figure 6: An example of the two-dimensional unbinned extraction of $S_{5}$ in the kinematic variables $q^{2}$ and $m_{K\pi}$. The pseudo-experiment data set has been generated with a relativistic Breit-Wigner in $m_{K\pi}$, which for this demonstration factorises with the angular distribution and $q^{2}$ (except for phase-space effects).
The zero-crossing points of the optimised observables that will be discussed from a theoretical point of view in the next section, may be readily ascertained with this method. An example of the extraction of the crossing points for $P_{2}$ and $P_{5}^{\prime}$ is shown in Fig. 7. As only the value of $q^{2}$ where these observables are zero is of interest, the crossing points are most easily discerned from the $\tilde{S}_{i}$ observables that come straight out of the $knn$ algorithm. These are shown for a single example pseudo-experiment; clearly there is discrimination between the SM and NP1 scenarios. Nevertheless, for a given single measurement the estimated uncertainty of the observable at a particular point in $q^{2}$ depends on the number neighbours and smoothing. Therefore, the uncertainty on the value of the crossing point must be be assessed with an ensemble of pseudo-experiments. The distribution of measured crossing points in $P_{2}$ and $P_{5}^{\prime}$, resulting from $\sim 10^{3}$ such experiments, is shown in Fig. 8. The expected uncertainty on these values is the width of these distributions. For $P_{2}$ and $P_{5}^{\prime}$, separation between SM and NP is clear; although with these sample sizes it is not possible to distinguish between the two NP scenarios. Importantly, the method is unbiased i.e. the centres of the distributions align with the true values used to generate the pseudo-experiments.
Figure 7: Extraction of the zero-crossing points of the observables (top) $S_{6}^{s}$ and (bottom) $S_{5}$, corresponding to the zeros of $P_{2}$ and $P_{5}^{\prime}$ respectively, in a single pseudo-experiment. For each, an example of the (blue) SM and (orange) a NP model is shown. The vertical lines denote the obtained central values and intervals for where the observables cross zero.
Figure 8: Expected precision for the extraction of the zero-crossing points. Shown are the variables (top) $S_{6}^{s}$ and (bottom) $S_{5}$, corresponding to $P_{2}$ and $P_{5}^{\prime}$, respectively. Comparisons are made with (blue) the SM predictions, (orange) NP scenario 1 and (green) NP scenario 2. The distributions of the pseudo-experiment results are the translucent histograms, with the predicted values given by the solid lines.
The geometrical bounds discussed in the next section may also be extracted in a slightly altered form. Taking the square of the bounds, to avoid potentially troublesome square roots, leads to the distributions in Fig. 9. It is seen that, in these single pseudo-experiments, the uncertainty bands of the measurements may violate the geometric bounds due to the limited statistics of the data. This is not unexpected. The true values of the relations get very close to the bounds and the experimental measurements treat each angular observable as a statistically independent quantity, that has no prior for the physical limits.
Figure 9: The square of the geometric bounds, as extracted from a single pseudo-experiment.
IV New geometrical bounds, analysis of zeroes and extracting Wilson coefficients
For the first time, we have model-independent access to the shape in the relevant $q^{2}$-regions of all the observables. This opens a new window of sensitivity to the Wilson coefficients and the possibility to implement robustness tests on data. In this section, we explore from the theoretical side, possible ways to extract information from the knowledge of the shape of the observables. In particular, we discuss new geometrical bounds, a well-known relation among the observables and we obtain the zeroes of three of the optimised observables. We also explore the robustness of the bounds and the symmetry relation in the presence of a NP scalar contribution.
IV.1 Description of new geometrical bounds and non-trivial relation with and without a scalar
In the absence of scalar contributions, as was assumed in the LHCb analysis of Ref. Aaij et al. (2024a), and neglecting lepton masses, there are 11 coefficients for the P-wave part of the angular distribution. Two of them are trivially related (see Ref. Egede et al. (2010)) but there is also one non-trivial relation. This implies that there is a maximum of eight independent observables that correspond to the eight degrees of freedom mentioned in subsection II.2. When taking into account the lepton mass terms, there are 11 coefficients and one relation is maintained (the non-trivial one, up to factors of $\beta$) and hence two extra degrees of freedom appear. Finally, in the presence of a scalar contribution and including lepton mass terms, there are eight complex amplitudes, $A_{\perp,\|,0}^{L,R}$ and $A_{t,S}$ ($n_{A}=8$), twelve coefficients ($n_{c}=12)$ and in this case there is no non-trivial relation ($n_{rel}=0$) and hence twelve degrees of freedom. Notice that we are focusing in this paper on the P-wave, for the S-wave we refer the reader to Ref. Algueró et al. (2021).
The relations between the observables are a consequence of the symmetries of the angular distribution. For instance, in the massless case the relation
$2n_{A}-n_{sym}=n_{c}-n_{rel},$ (29)
implies that there must be four symmetries of the distribution ($n_{sym}=4$) because there are six complex amplitudes $A_{\perp,\|,0}^{L,R}$ ($n_{A}=6$), eleven coefficients ($n_{c}=11)$ and the three relations ($n_{rel}=3$) explained above. The symmetries can be expressed in a compact form
$n^{\prime}=Un,$ (30)
where $U$ is a unitary transformation.
At this point we will present two different type of tests that can be used in the context of unbinned analyses: i) geometrical tests associated with the structure of the observables; and ii) tests based on the non-trivial relation between the observables.
The bounds and the non-trivial relation are strictly defined for unbinned observables, even if they have previously been applied only on binned observables Matias and Serra (2014). Since the non-trivial relation has been already discussed in several different papers Egede et al. (2010); Matias et al. (2012); Matias and Serra (2014) we refer the reader to these references for further details. Our interest here is to make the first tests of the relation directly with the unbinned amplitudes and observables, and to explore the size of the breaking of the relation and these bounds in the presence of a scalar contribution, where in principle some of them may no longer hold.
Figure 10: Bounds on optimised observables using the amplitude parameterisation of Ref. Aaij et al. (2024a) in the two cases $q^{2}>0$ and $q^{2}<0$ discussed in subsection II.3.
IV.1.1 Geometrical tests associated with the structure of the observables
The geometrical tests rely on the description of the observables in terms of the $n_{i}$ vectors and lead to bounds on the maximum size of the observables. There are three such tests:
  • i)
    Geometrical test I: One can combine the normalised observable Matias et al. (2012)
    $P_{4}=\frac{{\rm Re}(n_{0}^{\dagger}n_{\|}+{\rm CP})}{\sqrt{(|n_{\|}|^{2}+{\rm CP})(|n_{0}|^{2}+{\rm CP})}},$ (31)
    where CP stands for the CP conjugate and the relation with the $P_{4}^{\prime}$ observable given by $P_{4}=P_{4}^{\prime}/\sqrt{1-P_{1}}$ together with
    $P_{6}=\frac{{\rm Im}(n_{0}^{\dagger}n_{\|}+{\rm CP})}{\sqrt{(|n_{\|}|^{2}+{\rm CP})(|n_{0}|^{2}}+{\rm CP})},$ (32)
    and its corresponding relation with the $P_{6}^{\prime}$ observable given by $P_{6}=\beta P_{6}^{\prime}/\sqrt{1-P_{1}}$ to obtain
    $\quad\quad|P_{4}+iP_{6}|^{2}\leq 1\quad{\rm or}\quad|P_{4}^{\prime}+i\beta P_{6}^{\prime}|^{2}\leq 1-P_{1}.$ (33)
    This turns into
    $|P_{4}^{\prime}|\leq\sqrt{1-P_{1}-\beta^{2}P_{6}^{\prime 2}},$ (34)
    and equivalently
    $|P_{6}^{\prime}|\leq\frac{1}{\beta}\sqrt{1-P_{1}-P_{4}^{\prime 2}}.$ (35)
  • ii)
    Geometrical test II: Also in the case of $P_{5}^{\prime}$ and $P_{8}^{\prime}$ following the same strategy and taking as a starting point
    $P_{5}=\frac{{\rm Re}(n_{0}^{\dagger}n_{\perp}+{\rm CP})}{\sqrt{(|n_{\perp}|^{2}+{\rm CP})(|n_{0}|^{2}+{\rm CP})}},$ (36)
    together with $P_{5}=\beta P_{5}^{\prime}/\sqrt{1+P_{1}}$ and also
    $P_{8}=\frac{{\rm Im}(n_{0}^{\dagger}n_{\perp}+{\rm CP})}{\sqrt{(|n_{\perp}|^{2}+{\rm CP})(|n_{0}|^{2}}+{\rm CP})},$ (37)
    together with $P_{8}=P_{8}^{\prime}/\sqrt{1+P_{1}}$ one obtains
    $\quad\quad|P_{5}+iP_{8}|^{2}\leq 1\quad{\rm or}\quad|\beta P_{5}^{\prime}+iP_{8}^{\prime}|^{2}\leq 1+P_{1},$ (38)
    which turns into
    $|P_{5}^{\prime}|\leq\frac{1}{\beta}\sqrt{1+P_{1}-P_{8}^{\prime 2}},$ (39)
    and equivalently
    $|P_{8}^{\prime}|\leq\sqrt{1+P_{1}-\beta^{2}P_{5}^{\prime 2}}.$ (40)
  • iii)
    Geometrical test III: Finally, a bound on $P_{2}$ and $P_{3}$ can be derived following a different strategy. The starting point here is
    $P_{2}=\frac{1}{\beta}\frac{{\rm Re}(n_{\perp}^{\dagger}n_{\|}+{\rm CP})}{(|n_{\perp}|^{2}+|n_{\|}|^{2}+{\rm CP})},$ (41)
    and
    $P_{3}=\frac{{\rm Im}(n_{\perp}^{\dagger}n_{\|}+{\rm CP})}{(|n_{\perp}|^{2}+|n_{\|}|^{2}+{\rm CP})}.$ (42)
Combining these expressions and using $(|n_{\perp}|-|n_{\|}|)^{2}\geq 0$ one finds
$|P_{2}+i\frac{1}{\beta}P_{3}|\leq\frac{1}{2\beta},$ (43)
which implies
$|P_{2}|\leq\frac{1}{2\beta}\sqrt{1-4P_{3}^{2}},$ (44)
or alternatively,
$|P_{3}|\leq\frac{1}{2}\sqrt{1-4\beta^{2}P_{2}^{2}}.$ (45)
These bounds must be respected both in the SM and in presence of NP, and therefore they are a model-independent consistency test of the validity of an experimental analysis.
It is important to notice when comparing to data that the LHCb collaboration use a different convention for the observables than the original definition in Ref. Descotes-Genon et al. (2013b). These differing conventions can be interchanged using,
$P_{1}^{\rm LHCb}=P_{1},P_{2,3}^{\rm LHCb}=-P_{2,3},P_{4}^{\prime\rm LHCb}=-1/2P_{4}^{\prime},$
$P_{5,6}^{\prime\rm LHCb}=P_{5,6}^{\prime},P_{8}^{\prime\rm LHCb}=-1/2P_{8}^{\prime}.$ (46)
The previous bounds and the relations discussed below are expressed in the original basis.
In Fig. 10 the bounds are applied in the case of the amplitude analysis of Ref. Aaij et al. (2024a) discussed above. In an amplitude analysis the bounds must be fulfilled by construction and they therefore serve as a cross-check of the experimental analysis procedure. As the bound will continue to be respected in any amplitude analysis, even in the presence of a NP scalar, the bound cannot be used to detect any scalar NP contribution.
Fig. 10 illustrates that the bounds all become systematically stronger for large values of $q^{2}$. This is particularly noticeable in the case of $P_{4}^{\prime}$, where the bound is quasi-saturated at $q^{2}=8$ GeV2. The $P_{2}$ observable also saturates the bound around 2 GeV2, where the observable hits its maximum value of $P_{2}\sim 1/2$.
In contrast to the situation with amplitudes, measurements of observables - such as those proposed in section III - can violate some bounds in the presence of a scalar contribution. Our next point is to explore the extent of this violation.
In the presence of a scalar contribution, the bounds on $P_{5}^{\prime}$ and $P_{6}^{\prime}$ are modified and become
$|P_{5}^{\prime}+\omega_{r}|\leq\frac{1}{\beta}\sqrt{1+P_{1}-P_{8}^{\prime 2}},$ (47)
and
$|P_{6}^{\prime}+\omega_{i}|\leq\frac{1}{\beta}\sqrt{1-P_{1}-P_{4}^{\prime 2}},$ (48)
where
$\omega_{r}=+\frac{1}{N}\sqrt{2}\frac{m_{\ell}}{\beta\sqrt{q^{2}}}[{\rm Re}(A_{\|}^{L}A_{S}^{*}+A_{\|}^{R*}A_{S})+{\rm CP}],$ (49)
and
$\omega_{i}=+\frac{1}{N}\sqrt{2}\frac{m_{\ell}}{\beta\sqrt{q^{2}}}[{\rm Im}(A_{\perp}^{L}A_{S}^{*}-A_{\perp}^{R*}A_{S})+{\rm CP}],$ (50)
with
$N=\sqrt{(|n_{\perp}|^{2}+|n_{\|}|^{2}+{\rm CP})(|n_{0}|^{2}+{\rm CP})}.$ (51)
The terms $\omega_{r,i}$ contain the scalar contributions. If the signs of $P_{5}^{\prime}$ and $\omega_{r}$ are the same the bound in Eq. 39 may be violated in an analysis measuring observables, indicating the presence of a scalar NP contribution according to Eq. 47. Similarly, if the signs of $P_{6}^{\prime}$ and $\omega_{i}$ are the same, the bound in Eq. 35 may also be violated according to Eq. 48. All other bounds are unaffected by the presence of a scalar.
In the following, we quantify how large the breaking of the bounds and of the non-trivial relation can be under the assumption of a scalar NP contribution with all other amplitudes SM-like. The goal is to find out if such a NP contribution is detectable or, on the contrary, if the bounds are very robust and cannot be broken in a detectable way.
IV.1.2 Tests based on the non-trivial relation between the observables
Another consistency test of an analysis can be made using the non-trivial relation between the coefficients of the angular distribution. Under the assumption of massive leptons, this relation is given Egede et al. (2010)
$\!\!J_{2c}=\,\frac{\beta_{\ell}^{2}J_{6s}(J_{4}J_{5}+J_{7}J_{8})+J_{9}(\beta_{\ell}^{2}J_{5}J_{7}-4J_{4}J_{8})}{4J_{2s}^{2}-\left(J_{3}^{2}+\beta_{\ell}^{2}J_{6s}^{2}/4+J_{9}^{2}\right)}\!\!\,\,$ (52)
$-\,\frac{(2J_{2s}+J_{3})\left(4J_{4}^{2}+\beta_{\ell}^{2}J_{7}^{2}\right)+(2J_{2s}-J_{3})\left(\beta_{\ell}^{2}J_{5}^{2}+4J_{8}^{2}\right)}{8J_{2s}^{2}-\left(2J_{3}^{2}+\beta_{\ell}^{2}J_{6s}^{2}/2+2J_{9}^{2}\right)}.\quad\quad$
This relation has been already discussed and presented in Refs. Matias et al. (2012); Egede et al. (2010); Matias and Serra (2014) in the case of binned observables, where it is necessary to add a systematic uncertainty to handle the variation of the observables within each bin. However, this is the first time it can be tested directly on unbinned observables.
A non-trivial relation can also be obtained for the CP-conjugate coefficients, which leads to a formally identical combination.
At this point there are two options: 1) to test directly the relation in Eq. 52, which is given in terms of $J_{i}$; or 2) to re-express it in terms of the $P_{i}$ and $P_{i}^{CP}$ observables, as shown below. The latter is implemented by inverting the definitions in Eq. 12.
There is a minor difference between the two options, because in the first case a breaking of Eq. 52 can only be due to an internal problem of the experimental analysis, or to a large new scalar contribution. In the second case, if we use the expression in terms of $P_{i}$ and $P_{i}^{CP}$ and then assume that all $P_{i}^{CP}$ are negligible, we introduce a third possible source of breaking if there exists large NP weak phases in Nature. In any case, this relation is always a very efficient cross-check of an experimental analysis.
We have explicitly checked that, up to machine precision, the relation of Eq. 52 using the $J$’s obtained from the fit of Ref. Aaij et al. (2024a) is fulfilled, considering either the $q^{2}<0$ or $q^{2}>0$ cases. Since the $J_{i}$ are obtained from an amplitude analysis where no scalar contribution is introduced, the relation in Eq. 52 must, by construction, be fulfilled. However, this implies nothing about the presence or absence of a scalar amplitude.
In order to express Eq. 52 in terms of optimised observables we use the definitions of these observables in Eq. 12. One can then write down one relation for $\bar{P_{i}}=P_{i}+P_{i}^{\rm CP}$ and one for $\hat{P_{i}}=P_{i}-P_{i}^{\rm CP}$. Taking the example of $P_{2}$, we can write an equation for $\bar{P}_{2}$ and one for $\hat{P}_{2}$ as a function of all the other observables. The corresponding equations are Matias and Serra (2014)
${P_{2}+P_{2}^{\rm CP}}=+\frac{1}{2{\bar{k}_{1}}}\bigg{[}({\bar{P}_{4}^{\prime}}{\bar{P}_{5}^{\prime}}+\delta_{1})+\frac{1}{\beta}\sqrt{(-1+{\bar{P}_{1}}+{\bar{P}_{4}^{\prime 2}})(-1-{\bar{P}_{1}}+\beta^{2}{\bar{P}_{5}^{\prime 2}})+\delta_{2}+\delta_{3}{\bar{P}_{1}}+\delta_{4}{\bar{P}_{1}}^{2}}\bigg{]},$ (53)
and
${P_{2}-P_{2}^{\rm CP}}=+\frac{1}{2{\hat{k}_{1}}}\bigg{[}({\hat{P}_{4}^{\prime}}{\hat{P}_{5}^{\prime}}+\hat{\delta}_{1})+\frac{1}{\beta}\sqrt{(-1+{\hat{P}_{1}}+{\hat{P}_{4}^{\prime 2}})(-1-{\hat{P}_{1}}+\beta^{2}{\hat{P}_{5}^{\prime 2}})+\hat{\delta}_{2}+\hat{\delta}_{3}{\bar{P}_{1}}+\hat{\delta}_{4}{\hat{P}_{1}}^{2}}\bigg{]},$ (54)
where in these expressions $P_{2}$ stands for the value of the observable obtained from the relation, which should be exactly the same as the measured observable $P_{2}$ in the absence of a scalar contribution. The definitions of all $\bar{k}_{1}$ and ${\delta}_{i}$ can be found in Ref. Matias and Serra (2014). The corresponding $\hat{\delta}_{i}$ coefficients are simply obtained from the $\delta$’s in Ref. Matias and Serra (2014) by changing “bars” to “hats”. It is important to emphasise that this relation mixes $P_{i}$ and $P_{i}^{\rm CP}$ observables. In the SM, however, the $P_{i}^{\rm CP}$ are negligible and this implies that, to an excellent approximation, the relation can be simplified by taking all $P_{i}^{\rm CP}\to 0$. The simplified relation with $P_{i}^{\rm CP}\to 0$ can be broken in the presence of a large weak new physics phase that induces a non negligible $P_{i}^{\rm CP}$, or in the presence of a new scalar contribution. However, in this paper we will not consider further the possibility of a new weak phase, and instead will focus on analysing the breaking of this relation due to the presence of a NP scalar.
A simplified version of Eqs. 53 and 54, fulfilled to a high accuracy, is obtained in the approximation $P_{i}^{\rm CP}\to 0$, $P_{3}$ and $P_{6,8}^{\prime}\to 0$, which is equivalent to assuming no large NP weak and strong phases. This simplified version is given by
$P_{2}=\frac{1}{2}[P_{4}^{\prime}P_{5}^{\prime}+\Omega],$ (55)
where $\Omega=\frac{1}{\beta}\sqrt{(-1+P_{1}+{P_{4}^{\prime}}^{2})(-1-P_{1}+\beta^{2}{P_{5}^{\prime}}^{2}})$. For illustration, we show in Fig. 11 the expression in Eq. 55 rewritten in quadratic form using the results from one pseudo-experiment.
As discussed at the beginning of this subsection IV.1, in the presence of a NP scalar the non-trivial relation between the observables is broken. Here we estimate, for the first time, the amount of breaking of the relation due to the presence of a scalar by combining theory inputs with data on the branching fraction of ${B_{s}\to\mu^{+}\mu^{-}}$. If we allow for the presence of a scalar, Eq. 53 and Eq. 54 can be generalised by introducing the following substitutions in Eq. 53
$\bar{P_{5}^{\prime}}\to\bar{P_{5}^{\prime}}+\bar{\omega}_{r},\quad\quad\bar{P_{6}^{\prime}}\to\bar{P_{6}^{\prime}}+\bar{\omega}_{i},$ (56)
and in Eq. 54
$\hat{P_{5}^{\prime}}\to\hat{P_{5}^{\prime}}+{\hat{\omega}_{r}},\quad\quad\hat{P_{6}^{\prime}}\to\hat{P_{6}^{\prime}}+\hat{\omega}_{i},$ (57)
where $\hat{\omega}_{r,i}=\omega_{r,i}-\omega^{\rm CP}_{r,i}$ and $\bar{\omega}_{r,i}=\omega_{r,i}+\omega^{\rm CP}_{r,i}$ with
$\omega_{r}^{\rm CP}=+\frac{1}{N}\sqrt{2}\frac{m_{\ell}}{\beta\sqrt{q^{2}}}[{\rm Re}(A_{\|}^{L}A_{S}^{*}+A_{\|}^{R*}A_{S})-{\rm CP}],$ (58)
and
$\omega_{i}^{\rm CP}=+\frac{1}{N}\sqrt{2}\frac{m_{\ell}}{\beta\sqrt{q^{2}}}[{\rm Im}(A_{\perp}^{L}A_{S}^{*}-A_{\perp}^{R*}A_{S})-{\rm CP}],$ (59)
are the CP counterparts of the $\omega_{r,i}$ defined in Eq. 49 and Eq. 50, with $N$ given by Eq. 51. By expanding the terms in $\omega$, one can understand the maximal amount of breaking expected in the presence of a NP scalar.
Figure 11: Plot of the non-trivial relation rewritten as $W=P_{1}^{2}+P_{4}^{\prime 2}+\beta^{2}(4P_{2}^{2}-4P_{2}P_{4}^{\prime}P_{5}^{\prime}+P_{5}^{\prime 2})+P_{1}(P_{4}^{\prime 2}-\beta^{2}P_{5}^{\prime 2})$, where $W=1$ for all $q^{2}$. Only one pseudo-experiment is used for illustrative purposes.
[…]
Figure 13: Plot of the position of the zero of $P_{2}$ and $P_{4,5}^{\prime}$ versus $C_{9}^{\rm NP}$. The dashed line corresponds to the approximate expression of Eq. 68.
As an illustration, assuming neither large NP weak phases nor large strong phases, one can take the simplest form of the relation and expand in $\omega_{r}$. Assuming a small amount of breaking in $\omega_{r}$, one finds
$P_{2}=\frac{1}{2}\left[P_{4}^{\prime}P_{5}^{\prime}+\Omega\right]+\Delta,$ (60)
and
$\Delta=\frac{\omega_{r}}{2}\left(P_{4}^{\prime}-\beta P_{5}^{\prime}\sqrt{\frac{-1+P_{1}+{P_{4}^{\prime}}^{2}}{-1-P_{1}+\beta^{2}{P_{5}^{\prime}}^{2}}}\right).$ (61)
[…]
The small size of $\omega_{r,i}$, and therefore of $\Delta$ in Eq. (61), means that one cannot discern if there is a scalar contribution to $B\to K^{*}\mu^{+}\mu^{-}$ from the breaking of the relations and the bounds. This is consistent with previous determinations of these contributions within WET analyses of $b\to s\ell^{+}\ell^{-}$ data Hurth et al. (2023); Beaujean et al. (2015b); Bobeth et al. (2013).
In summary, we have found that the bounds and the non-trivial relation assuming the absence of large new weak phases should be fulfilled to a high-accuracy, making them a robust internal cross-check of any experimental analysis.
IV.2 Zero-crossing points of the optimised observables
Knowledge of the shape of the observables $P_{1,2}$ and $P_{4,5}^{\prime}$ provides information on zero-crossing points, which can be used to determine the Wilson coefficients.
The best known zero-crossing point is that of the forward-backward asymmetry, which is equivalent to the zero-crossing point of the $P_{2}$ observable. In the limit of the initial hadron being heavy and the final meson having a large energy, i.e., at small $q^{2}$, the form factors are related by
$\frac{T_{1}(q^{2})}{V(q^{2})}=\frac{m_{B}^{2}}{(m_{B}+m_{K^{*}})^{2}}\frac{T_{2}(q^{2})}{A_{1}(q^{2})}$ (62)
and using that $T_{1}(q^{2})/V(q^{2})=m_{B}/(m_{B}+m_{K^{*}})$ in this same limit, all form factor dependence drops out and one gets,
${q_{0}^{2}}^{LO}\Big{\rvert}_{P_{2}}=-2m_{b}m_{B}\frac{{\cal C}_{7}^{\rm eff}}{{\rm Re}{\cal C}_{9}^{\rm eff}(q_{0}^{2})}.$ (63)
This is the well known expression of the position of the zero-crossing point at leading order Ali et al. (1995). However, in this case, the next-to-leading order correction is substantial and moves the crossing-point in the SM from $\simeq 3$ GeV2 to $q_{0}^{2}=3.60\pm 0.40$ GeV2. Unfortunately, the explicit expression for the next-to-leading order zero-crossing point is a complicated function of form factors, hadronic nuisance parameters, as well as the Wilson coefficients. The global analysis of Ref. Algueró et al. (2023) points to NP only in the Wilson coefficient ${\cal C}_{9}$, with marginal contributions to the other Wilson coefficients. We therefore predict, at next-to-leading order, the theoretical position of the zero of $P_{2}$ as a function of ${\cal C}_{9}^{\rm NP}$, as shown in Fig. 13. A strong dependence of the zero-crossing points of $P_{2}$ and $P_{5}^{\prime}$ on ${\cal C}_{9}^{\rm NP}$ is observed but there is a marginal dependence on the crossing point of $P_{4}^{\prime}$. We have not added the corresponding zero-crossing point of $P_{1}$, because it is relatively flat and therefore has significant uncertainties. Nonetheless, we explain below a further interesting feature of the $P_{1}$ zero-crossing point.
Using the non-trivial symmetry relation we can find an alternative way of cross-checking the position of the zero of $P_{2}$ in the SM and in the presence of NP. It consists of finding where the following function is zero
$X_{P_{2}}=P_{4}^{\prime}P_{5}^{\prime}+\Omega.$ (64)
This is true up to corrections of ${\cal O}(P_{i}^{CP},P_{3},P_{6,8}^{\prime})$, i.e., the same approximation as in Eq. 55.
The zero-crossing point of $P_{5}^{\prime}$ in the SM is $q_{0}^{2}=2.16\pm 0.29$ GeV2. Again here the exact expression is rather cumbersome, so instead of writing it explicitly we present the exact zero-crossing point as a function of ${\cal C}_{9}^{\rm NP}$ in Fig. 13. As in the case of $P_{2}$, there also exists for $P_{5}^{\prime}$ an alternative way to determine its zero using the expression,
$X_{P_{5}^{\prime}}=-P_{2}+\frac{1}{2\beta}\sqrt{(-1+P_{1}+P_{4}^{\prime 2})(-1-P_{1})}.$ (65)
An interesting property is found when comparing the zeroes of $P_{5}^{\prime}$ and $P_{1}$. To a high accuracy both observables have the same zero-crossing point in the SM, i.e.
${q_{0}^{2}}\Big{\rvert}_{P_{5}^{\prime}}^{\rm SM}\simeq q_{0}^{2}\Big{\rvert}_{P_{1}}^{\rm SM}.$ (66)
[…]
$\mathcal{C}_{S}$ ($\mathrm{GeV}^{-1}$) 0.000 [-0.0084, 0.0084] [-0.0129, 0.0129]
[…]
On the contrary, in the presence of NP in ${\cal C}_{9}^{\rm NP}$ the zeroes of $P_{1}$ and $P_{5}^{\prime}$ are affected differently. Specifically, when adding a ${\cal C}_{9}^{\rm NP}=-1.17$, the central value of the zero-crossing point of $P_{1}$ moves marginally by $+0.10$ GeV2, while the crossing point of $P_{5}^{\prime}$ moves by $+0.85$ GeV2. This means that, independently of the exact position of the zeroes of $P_{1}$ and $P_{5}^{\prime}$, if these crossing points do not coincide it indicates a NP contribution to ${\cal C}_{9}$. Unfortunately when adding uncertainties to the zero-crossing point of $P_{1}$, the mild dependence on $C_{9}$ makes this difficult to exploit in experiments.
A byproduct of this is that at this common zero in the SM, using the simplified form of the relation between the observables, one finds
${P_{4}^{\prime}}^{2}+4\beta^{2}{P_{2}^{2}}\Big{\rvert}_{P_{5}^{\prime},P_{1}}^{\rm SM}\simeq 1.$ (67)
Finally, for the case of $P_{4}^{\prime}$, given that the position of its zero is even closer to the limit of large energy for the $K^{*}$ (small $q^{2}$) the expressions in the limit of large $E_{K^{*}}$ work accurately. At next-to-leading order, the zero-crossing point of $P_{4}^{\prime}$ in the SM is found at $q_{0}^{2}=1.57\pm 0.25$ GeV2 and in the presence of a NP contribution $C_{9}^{\rm NP}\sim-1$ it becomes even closer to $q^{2}\simeq 0$ GeV2.
In this limit, the position of the zero-crossing point for $P_{4}^{\prime}$ in the SM is found to be
${q_{0}^{2}}\Big{\rvert}_{P_{4}^{\prime}}=-2m_{b}{\cal C}_{7}^{\rm eff}\frac{(2m_{b}{\cal C}_{7}^{\rm eff}+m_{B}{\rm Re}{\cal C}_{9}^{\rm eff}(q_{0}^{2}))}{({{\cal C}_{10}^{\rm eff}}^{2}+{|{\cal C}_{9}^{\rm eff}(q_{0}^{2})|}^{2})}$ (68)
in excellent agreement with the next-to-leading order SM prediction, as can be seen in Fig. 13.
The $C_{9}$ dependence of the zero of $P_{4}^{\prime}$ is small, such that the ability of experimental data to distinguish SM and NP is limited. However, as shown in Figs. 7 and 8, the new experimental method presented in this paper will have excellent ability to resolve expected NP effects in the observables $P_{2}$ and $P_{5}^{\prime}$.
V Conclusions and Outlook
In this paper we present a model-independent unbinned method, based on a moments approach, to extract angular observables in the four-body decay mode $B\to K^{*}(\to K^{+}\pi^{-})\mu^{+}\mu^{-}$. This method is an alternative to the present approach of putting the data into fixed bins of $q^{2}$ and maintains the model independence of the observables. It will also allow for the first studies of the $m_{K\pi}$ dependence of the observables, shedding new light on how the P- and S-wave observables depend on the $m_{K\pi}$ kinematic variable. The method is in contrast to different amplitude analyses that have been presented, which are model-dependent. In the case of experimental amplitude analyses, we define a procedure to test the validity of a theoretical model for the amplitudes. As an illustration, we apply this test to the fit of Ref. Gubernari et al. (2021), fixing the theory parameters at $q^{2}<0$. We show that one of the conditions is not fulfilled and suggest several possible reasons for this.
We provide the SM predictions for all the unbinned optimised observables and derive a set of new geometrical bounds on these observables. We explore the impact of a NP scalar on these bounds and on a non-trivial relation between the unbinned observables, considering the constraint from the branching fraction of the decay $B_{s}\to\mu^{+}\mu^{-}$. We find that the impact of scalar NP is negligible and therefore the bounds and the relation are robust, and can be used as a cross-check of an experimental analysis procedure.
We also provide next-to-leading order predictions for the zero crossing-points of the optimised observables $P_{2}$ and $P_{4,5}^{\prime}$ as a function of $C_{9}^{\rm NP}$, and show that they can be extracted with the proposed moments approach.
In a forthcoming paper Capdevila et al. (in preparation), we will investigate the precision with which the Wilson coefficients can be extracted with the unbinned moments approach and compare to a traditional binned fit.
Acknowledgements.
The authors acknowledge Konstantinos Petridis and Michael McCann for enlightening discussions. B.C. would like to thank the Phenomenology Group at DAMTP for useful discussions at the early stages of this work. J.M. gratefully acknowledges the financial support from ICREA under the ICREA Academia programme 2018 and to AGAUR under the Icrea Academia programme 2024 and from Departament de Recerca i Universitats de la Generalitat de Catalunya. J.M. also received financial support from the Spanish Ministry of Science, Innovation and Universities (project PID2020-112965GB-I00) and from the IPPP Diva Award 2024. The work of B.C. is supported by the Margarita Salas postdoctoral program funded by the European Union-NextGenerationEU and the La Caixa Junior Leader fellowship from the “la Caixa” Foundation (ID 100010434, fellowship code LCF/BQ/PI24/12040024). MN acknowledges the financial support by the Spanish Government (Agencia Estatal de Investigación MCIN/AEI/10.13039/501100011033) and the European Union NextGenerationEU/PRTR through the “Juan de la Cierva” program (Grant No. JDC2022-048787-I), Grants No. PID2020-114473GB-I00 and No. PID2023-146220NB-I00. M.P. and M.S. acknowledge financial support from the UK Science and Technology Facilities Council (STFC).
Appendix A SM UNBINNED PREDICTIONS OF NON-OPTIMISED OBSERVABLES
For completeness, we provide also the SM predictions of the unbinned non-optimised observables.
Figure 14: SM prediction for the unbinned non-optimised observables $S_{3,4,5,7,8,9}$, $A_{\rm FB}$ and the branching fraction.
Appendix B CONDITIONS
The rest of the conditions that exhibit a perfect agreement are presented in Fig. 15.
Figure 15: Illustration of conditions 1, 2, 3, 5, 6 and 7. These use the same convention as in Fig. 3.
Appendix C LEPTON MASS MODIFIED DISTRIBUTION
In the theory convention the four-body angular distribution is written:
$\frac{d^{4}\Gamma}{dq^{2}\,d\!\cos\theta_{K}\,d\!\cos\theta_{l}\,d\phi}=\frac{9}{32\pi}\bigg{[}J_{1s}\sin^{2}\theta_{K}+J_{1c}\cos^{2}\theta_{K}+(J_{2s}\sin^{2}\theta_{K}+J_{2c}\cos^{2}\theta_{K})\cos 2\theta_{l}$ (69)
$\hskip 76.82234pt+J_{3}\sin^{2}\theta_{K}\sin^{2}\theta_{l}\cos 2\phi+J_{4}\sin 2\theta_{K}\sin 2\theta_{l}\cos\phi+J_{5}\sin 2\theta_{K}\sin\theta_{l}\cos\phi$
$\hskip 76.82234pt+(J_{6s}\sin^{2}\theta_{K}+{J_{6c}\cos^{2}\theta_{K}})\cos\theta_{l}+J_{7}\sin 2\theta_{K}\sin\theta_{l}\sin\phi+J_{8}\sin 2\theta_{K}\sin 2\theta_{l}\sin\phi$
$\hskip 76.82234pt+J_{9}\sin^{2}\theta_{K}\sin^{2}\theta_{l}\sin 2\phi\bigg{]}\,X+S$
where S stands for the S-wave part of the amplitude (See Refs.Descotes-Genon et al. (2016, 2013b); Matias et al. (2012) for definitions).
There is a way to define a simpler distribution that contains only the $J$’s required to define the $P_{i}$ observables and $F_{L}$ and that avoids the conflicting terms $J_{1s}$ and $J_{1c}$ that are proportional to $F_{T}$ and $F_{L}$, respectively, in the massless limit, but that they contain corrections in the massive case.
The following difference of two distributions is a function of only the terms that are needed to define the relevant observables:
$d\hat{\Gamma}=d\Gamma(\theta_{K},\theta_{\ell},\phi)-d\Gamma(\theta_{K},\theta_{\ell}+\frac{\pi}{2},\phi)$ (70)
where
$d\hat{\Gamma}=$ $\frac{9}{32\pi}\bigg{[}2(J_{2s}\sin^{2}\theta_{K}+J_{2c}\cos^{2}\theta_{K})\cos 2\theta_{l}$ (71)
$-J_{3}\sin^{2}\theta_{K}\cos 2\theta_{l}\cos 2\phi+2J_{4}\sin 2\theta_{K}\sin 2\theta_{l}\cos\phi+J_{5}\sin 2\theta_{K}(\sin\theta_{l}-\cos\theta_{l})\cos\phi$
$+(J_{6s}\sin^{2}\theta_{K}+{J_{6c}\cos^{2}\theta_{K}})(\sin\theta_{l}+\cos\theta_{l})+J_{7}\sin 2\theta_{K}\sin\phi(\sin\theta_{l}-\cos\theta_{l})+2J_{8}\sin 2\theta_{K}\sin 2\theta_{l}\sin\phi$
$-J_{9}\sin^{2}\theta_{K}\cos 2\theta_{l}\sin 2\phi\bigg{]}\,X+\hat{S}$
where $\hat{S}$ stands for the transformed S-wave under the same condition. This new distribution could be useful to avoid problems of precision associated to the extraction of observables when the lepton mass is taken into account.
This difference of distributions does not help with the extraction of the observables from the method of moments when the lepton mass is not neglected, as the degeneracy amongst $S_{2}^{s}$, $S_{2}^{c}$ and $\tilde{S}_{2a}^{c}$ remains. However, for a fit of the data that includes the $m_{K\pi}$ dependence, as in the current generation of analyses, there should be an advantage from this strategy. This is due to the reduction of the degrees of freedom in such a fit.
Appendix D DISENTANGLING S-WAVE CONTRIBUTIONS AND MASSIVE LEPTONS
To demonstrate the treatment of data with an S-wave component, the pseudo-experiments of the main text have been augmented with a simple S-wave model. The S-wave contribution relative to the P-wave is set to be 0 at the kinematic endpoints ($q^{2}=0\,\text{GeV}^{2}$ and $q^{2}=19\,\text{GeV}^{2}$) and $\sim 10\%$ at $q^{2}=9\,\text{GeV}^{2}$. Here, the interference between P- and S-wave has been neglected in the generation of the pseudo-data for the sake of simplicity. The P-wave events are generated with a relativistic Breit Wigner in $m_{K\pi}$ and the S-wave with a LASS parametrisation.
In the case that lepton masses are non-negligible one must extract $S_{1}^{c}$, $S_{1}^{s}$, $S_{2}^{c}$ and $S_{2}^{s}$ as they are not trivially related to each other. The respective angular functions of these observables are not orthogonal, such that they can only be obtained from a combination of the moments. In the absence of S-wave contributions, the angular observables are obtained from the moments as per Eq. III.
Assuming the leptons are massless (and ignoring tensor contributions), one may use the relations Altmannshofer et al. (2009)
$J_{1}^{c}=-J_{2}^{c},$ $J_{1}^{s}=3J_{2}^{s},$ $J_{1a}^{c}=-J_{2a}^{c},$ (72)
in which case determining the observables from the moments is trivial, as in Ref. Aaij et al. (2016a). The analyst can pick a middle ground and not use all the relations of massless leptons, to avoid the most egregious biases from lepton masses, as long as there is a unique solution for the observables from the moments.
By taking the leptons as massless one can study the $m_{K\pi}$ dependence of the observables for $q^{2}>\sim 4\,\text{GeV}^{2}$, which achieves several aims. The $m_{K\pi}$ dependence of the P-wave observables will be assessed for the first time, shedding new light on the narrow-width approximation for the $K^{\ast{}0}$. One can also extract new information about the S-wave observables, in particular if the commonly used functions (the LASS parametrisation or some isobar model) are accurate. It will also become clear if the amplitudes that describe the S-wave should be considered to be the same at low and high values of $m_{K\pi}$, as is currently assumed by the factorisation of $m_{K\pi}$ from the other kinematic variables in the calculation of the angular observables. The main text highlights the possibility of achieving a fully two-dimensional map in $m_{K\pi}$ and $q^{2}$ of the P-wave observables (without assuming factorisation) in Fig. 6.
Alternatively, the analyst may incorporate information from the kinematic variable $m_{K\pi}$. Considering $m_{K\pi}$, the differential decay rate may be written as
$\frac{{\rm d}\Gamma}{{\rm d}q^{2}{\rm d}\Omega{\rm d}m_{K\pi}}=$
$\frac{9}{32\pi}\left[\sum_{i=1s\to 9}J_{i}(q^{2})f_{i}(\Omega)|f_{P}(m_{K\pi}|q^{2})|^{2}\right]$
$+\frac{1}{4\pi}\left[\sum_{i=1ac,2ac}J_{i}(q^{2})f_{i}(\Omega)|f_{S}(m_{K\pi}|q^{2})|^{2}\right]$
$+\frac{1}{4\pi}\left[\sum_{i=1bc\to\tilde{S}_{8}}J_{i}(q^{2})f_{i}(\Omega)f_{s}(m_{K\pi}|q^{2})f_{P}(m_{K\pi}|q^{2})\right],$ (73)
where $\Omega$ represents the three decay angles. The first line describes the P-wave decay, the second is the S-wave and the third the interference between P- and S-wave. The terms $f_{P}(m_{K\pi}|q^{2})$ and $f_{S}(m_{K\pi}|q^{2})$ are complex analytic functions describing the P- and S-wave distributions in $m_{K\pi}$ for a given point in $q^{2}$. There is an implicit assumption that $m_{K\pi}$ factorises from $\Omega$. This should be checked at high $q^{2}$ using the massless lepton assumption to determine if the angular observables are functions of $m_{K\pi}$ as well as $q^{2}$.
Figure 16: The determination of (top) $\tilde{S}_{1a}^{c}$ and (bottom) $\tilde{S}_{2a}^{c}$ when including the $m_{K\pi}$ description of the calculation of the moments. Shown here is the result of a single pseudo-experiment with the same S-wave amplitudes. The difference between SM and NP are due to the different predicted branching fractions of the P-wave components of the decay.
Figure 17: The determination of (green band) $\tilde{S}_{5}$ and (blue band) $S_{5}$ with the moments, compared with (blue line) the true predicted value of $S_{5}$ as from Table 1. The need to renormalise the observables that have been extracted from the moments to account for S-wave contributions is clear.
If the angular observables are not functions of $m_{K\pi}$, then $f_{S}(m_{K\pi}|q^{2})$ and $f_{P}(m_{K\pi}|q^{2})$ may be incorporated into the method of moments, for example as
$M_{1}^{s}=\int_{\rm min}^{\max}\int\frac{{\rm d}\Gamma}{{\rm d}\Omega{\rm d}m_{K\pi}}\sin^{2}\theta_{K}|f_{P}(m_{K\pi}|q^{2})|^{2}{\rm d}\Omega{\rm d}m_{K\pi},$ (74)
and analogously for the other moments. The angular functions are by and large orthogonal, so this is only pertinent to the coefficients $S_{1}^{s}$, $S_{1}^{c}$, $S_{s}^{s}$, $S_{s}^{c}$, $\tilde{S}_{1a}^{c}$ and $\tilde{S}_{2a}^{c}$. The precise form of $f_{P}(m_{K\pi}|q^{2})$ and $f_{S}(m_{K\pi}|q^{2})$ are up to the analyst and inevitably contribute as a source of systematic uncertainty in the experimental analysis. One must choose that the functions be normalised over the analysed $m_{K\pi}$ range
$\int_{\rm min}^{\rm max}|f_{P}(m_{K\pi}|q^{2})|^{2}{\rm d}m_{K\pi}=1,$
$\int_{\rm min}^{\rm max}|f_{S}(m_{K\pi}|q^{2})|^{2}{\rm d}m_{K\pi}=1.$ (75)
Defining the integrals
$A=\int_{\rm min}^{\rm max}|f_{P}(m_{K\pi}|q^{2})|^{4}{\rm d}m_{K\pi},$
$B=\int_{\rm min}^{\rm max}|f_{S}(m_{K\pi}|q^{2})|^{4}{\rm d}m_{K\pi},$
$C=\int_{\rm min}^{\rm max}|f_{P}(m_{K\pi}|q^{2})|^{2}|f_{S}(m_{K\pi}|q^{2})|^{2}{\rm d}m_{K\pi},$ (76)
one can write down the moments
$M_{1}^{s}=\frac{4}{9}C(3\tilde{S}_{1a}^{c}-\tilde{S}_{2a}^{c})+\frac{1}{10}A(3\tilde{S}_{1}^{c}+12\tilde{S}_{1}^{s}-\tilde{S}_{2}^{c}-4\tilde{S}_{2}^{s}),$
$M_{1}^{c}=\frac{2}{9}C(3\tilde{S}_{1a}^{c}-\tilde{S}_{2a}^{c})+\frac{1}{20}A(9\tilde{S}_{1}^{c}+6\tilde{S}_{1}^{s}-3\tilde{S}_{2}^{c}-2\tilde{S}_{2}^{s}),$
$M_{2}^{s}=\frac{1}{50}A(-5\tilde{S}_{1}^{c}-20\tilde{S}_{1}^{s}+7\tilde{S}_{2}^{c}+28\tilde{S}_{2}^{s})-\frac{4}{45}C(5\tilde{S}_{1a}^{c}-7\tilde{S}_{2a}^{c}),$
$M_{2}^{c}=\frac{1}{100}A(-15\tilde{S}_{1}^{c}-10\tilde{S}_{1}^{s}+21\tilde{S}_{2}^{c}+14\tilde{S}_{2}^{s})-\frac{2}{45}C(5\tilde{S}_{1a}^{c}-7\tilde{S}_{2a}^{c}),$
$M_{1a}^{c}=\frac{1}{60}(3C(-5\tilde{S}_{1}^{c}-10\tilde{S}_{1}^{s}+7\tilde{S}_{1}^{c}+14\tilde{S}_{2}^{s})-8B(5\tilde{S}_{1a}^{c}-7\tilde{S}_{2a}^{c})),$
$M_{2a}^{c}=\frac{1}{12}(3C(3\tilde{S}_{1}^{c}+6\tilde{S}_{1}^{s}-\tilde{S}_{2}^{c}-2\tilde{S}_{2}^{s})+24B\tilde{S}_{1a}^{c}-8B\tilde{S}_{2a}^{c}).$ (77)
One then finds
$\tilde{S}_{1}^{s}=\frac{1}{48A(C^{2}-AB)}\left[AC(20M_{2a}^{c}+28M_{1a}^{c})-(\frac{9}{5}AB-C^{2})(35M_{1}^{s}+25M_{2}^{s})+(\frac{3}{5}AB-C^{2})(70M_{1}^{c}+50M_{2}^{c})\right],$
$\tilde{S}_{2}^{s}=\frac{5}{48A(C^{2}-AB)}\left[AC(12M_{2a}^{c}+4M_{1a}^{c})-(\frac{9}{5}AB-C^{2})(5M_{1}^{s}-15M_{2}^{s})+(\frac{3}{5}AB-C^{2})(10M_{1}^{c}+30M_{2}^{c})\right],$
$\tilde{S}_{1}^{c}=\frac{1}{24A(C^{2}-AB)}\left[AC(10M_{2a}^{c}+14M_{1a}^{c})-(\frac{6}{5}AB-C^{2})(70M_{1}^{c}+50M_{2}^{c})+(\frac{3}{5}AB-C^{2})(35M_{1}^{s}+25M_{2}^{s})\right],$
$\tilde{S}_{2}^{c}=\frac{5}{24A(C^{2}-AB)}\left[AC(6M_{2a}^{c}+2M_{1a}^{c})-(\frac{6}{5}AB-C^{2})(10M_{1}^{c}+30M_{2}^{c})+(\frac{3}{5}AB-C^{2})(5M_{1}^{s}+15M_{2}^{s})\right],$
$\tilde{S}_{1a}^{c}=\frac{1}{32(C^{2}-AB)}\left[-15AM_{2a}^{c}-21AM_{1a}^{c}+C(21M_{1}^{s}+21M_{1}^{c}+15M_{2}^{s}+15M_{2}^{c})\right],$
$\tilde{S}_{2a}^{c}=\frac{15}{32(C^{2}-AB)}\left[-3AM_{2a}^{c}-AM_{1a}^{c}+C(M_{1}^{s}+M_{1}^{c}+3M_{2}^{s}+3M_{2}^{c})\right].$ (78)
Setting $C=0$, one recovers the P-wave only relations of Eq. III. This discussion is moot for the observables describing the interference between the P- and S-wave contributions, $\tilde{S}_{1b}^{c}$, $\tilde{S}_{2b}^{c}$, $\tilde{S}_{4}$, $\tilde{S}_{5}$, $\tilde{S}_{7}$ and $\tilde{S}_{8}$. These are already normalised by $\Gamma_{P}+\Gamma_{S}$ and are orthogonal to the other angular functions and so may be extracted directly from their respective moments.
Using the approach of including $m_{K\pi}$ in the moments calculations, as in Eq. 78 one has the correct extraction of $\tilde{S}_{1a}^{c}$ and $\tilde{S}_{2a}^{c}$, as shown in Fig. 16. Consequently one corrects the normalisation of the P-wave observables with a factor of $1/(1-2\tilde{S}_{1a}^{c}+\frac{2}{3}\tilde{S}_{2a}^{c})$. The effect of this renormalisation, and consequently its importance, is evident in Fig. 17.
It should be noted that concerns due to S-wave contributions only pertains to decays of the type $B\to V\ell^{+}\ell^{-}$, where there may be several partial waves contributing. When applying this method to $B\to P\ell^{+}\ell^{-}$ decays, such as $B^{+}\to K^{+}\mu^{+}\mu^{-}$ or $B^{+}\to\pi^{+}\mu^{+}\mu^{-}$, there are no other contributing partial waves and so the unbinned extraction of the angular coefficients is trivial.